OpenStudy (anonymous):

What would be the most logical first step to solve this quadratic equation? (x+5)^2=36 A.make a difference of squares so you can immediately factor . B.subtract 5 from both sides. C.take the square root of both sides. D.subtract 36 from both sides so you can immediately factor.

OpenStudy (anonymous):

would it be A

OpenStudy (anonymous):

weird. im a bit confused. one sec

OpenStudy (anonymous):

WAIT!!! C

OpenStudy (anonymous):

sure?

OpenStudy (anonymous):

A and D are pretty much the same...

OpenStudy (anonymous):

yup right ok ima go with c

OpenStudy (anonymous):

you could write \[x+5=6 \] or \[x+5=−6 \] as a first step and then solve

OpenStudy (radar):

since both sides are perfect squares, take the square root of both sides getting: x+5 = 6 then solve for x.

OpenStudy (radar):

Yes, you have already done that, very good ify214.

OpenStudy (mathmate):

A and D are essentially the same, but A is more logical and precise, IMHO.